# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v1

@@ 1 Introduction
Recent extensive research demonstrates that token compression substantially enhances inference efficiency, driving the continuous development of diverse compression strategies and sophisticated methodologies [43, 44, 27, 26, 45, 18, 28, 46, 47, 48, 25, 49, 50, 51]. However, the inherent heterogeneity of multimodal data means that redundancy manifests differently across modalities. Unlike textual prompts, where redundancy is primarily in syntactic or semantic, visual and auditory data exhibit unique structural properties. For instance, high-resolution images contain strong local correlations, while video streams feature extensive spatiotemporal redundancy across frames, and audio signals often contain extended segments of silence or stationary noise. Consequently, most existing methods focus on compressing one or two specific modalities.
@@
LongVU [94], TimeChat-Online [95], VidCom2 [96], DynTok [97], VideoChat-Flash [92],
Video-XL [98], LLaVA-Scissor [99] , leaf3, text width=44em] ] [   Attention-based
@@
iLLaVA [65], G-Search [79], FiCoCo [70] , leaf3, text width=44em] ] [   Query-based
@@ 2.3 Vision Transformer Token Compression
Visual token compression, originally pioneered in vision transformers (ViTs) [24, 146, 147, 148, 149, 150, 151], offers insights for addressing analogous challenges in MLLMs.
@@ 3.3 Attention-based Image-centric Compression
MLMMs w/ Token Compression
@@ 4 Video-centric Token Compression
To address this, current video LLMs commonly employ a 1 FPS sampling rate to reduce token counts. Furthermore, unlike the typical approach for single images where image is split into patches and fed into an encoder for detailed information, video processing often foregoes this detailed frame-level segmentation to keep token numbers manageable. Even with these strategies, the quantity of video tokens remains substantial. During model training and understanding, transformation-based methods, such as the pooling technique used in LLaVA-Video [33], are frequently employed to reduce tokens and aid the model’s comprehension of video content.
@@ 6.2.3 Enable One-Shot Long-Context Understanding
Limited by the inherent length of the context, MLLMs are unable to learn or comprehend real-world scenarios involving extremely long contexts, such as understanding entire code repositories or extended video and audio sequences. However, token compression significantly condenses and abstracts original information representations, making it possible for MLLMs to understand these long contexts in a single pass.
Traditional methods for handling long contexts in MLLMs, like FlashAttention [158, 159] or RingAttention [188], involve architectural changes to the model’s attention mechanism to directly accommodate longer sequences. While effective, these require fundamental model modifications. Token compression offers a different, often simpler, route. Instead of redesigning the model to fit more tokens, it focuses on making each token more powerful. By creating information-dense tokens, we pack more meaning into fewer pieces of data. This lets existing MLLM architectures process significantly longer conceptual contexts without major overhauls. It’s a more efficient and accessible way to achieve that crucial one-shot understanding of vast, complex real-world information.
@@ 6.4.2 Task-Specific Challenges
AuroraCap [44] Video Caption GPT-score Describe the video in detail.
@@ 6.6.1 Unified Multimodal Token Compression
Although different modalities exhibit unique redundancy patterns requiring specialized compression strategies, contemporary MLLMs predominantly focus on joint multimodal inference [10], whereas unimodal scenarios remain constrained by single-modality inputs. As established in Sections 3, 5, and 4, fundamental algorithmic principles (including transformation-based, similarity-based, attention-based, and query-based approaches) demonstrate transmodal applicability, indicating the viability of developing a unified multimodal token compression framework. Furthermore, leveraging cross-modal correlations represents a promising research direction for future token compression.

### v2

@@ 1 Introduction
Recent extensive research demonstrates that token compression substantially enhances inference efficiency, driving the continuous development of diverse compression strategies and sophisticated methodologies [44, 45, 28, 27, 46, 18, 29, 47, 48, 49, 26, 50, 51, 52, 25]. However, the inherent heterogeneity of multimodal data means that redundancy manifests differently across modalities. Unlike textual prompts, where redundancy is primarily in syntactic or semantic, visual and auditory data exhibit unique structural properties. For instance, high-resolution images contain strong local correlations, while video streams feature extensive spatiotemporal redundancy across frames, and audio signals often contain extended segments of silence or stationary noise. Consequently, most existing methods focus on compressing one or two specific modalities.
@@
LongVU [95], VidCom2 [96], DynTok [97], MovieChat [98], VideoChat-Flash [93], Video-XL [99],
TimeChat-Online [100], DTC [101], LLaVA-Scissor [102], Aim [103] , leaf3, text width=44em] ] [   Attention-based
@@
iLLaVA [66], G-Search [80], FiCoCo [71], Aim [103] , leaf3, text width=44em] ] [   Query-based
@@ 2.3 Vision Transformer Token Compression
Visual token compression, originally pioneered in vision transformers (ViTs) [24, 150, 151, 152, 153, 154, 155, 156], offers insights for addressing analogous challenges in MLLMs.
@@ 3.3 Attention-based Image-centric Compression
MLLMs w/ Token Compression
@@ 4 Video-centric Token Compression
To address this, current video LLMs commonly employ a 1 FPS sampling rate to reduce token counts. Furthermore, unlike the typical approach for single images, where the image is split into patches and fed into an encoder for detailed information, video processing often foregoes this detailed frame-level segmentation to keep token numbers manageable. Even with these strategies, the quantity of video tokens remains substantial. During model training and understanding, transformation-based methods, such as the pooling technique used in LLaVA-Video [34], are frequently employed to reduce tokens and aid the model’s comprehension of video content.
@@ 6.2.3 Enable One-Shot Long-Context Understanding
Limited by the inherent length of the context, MLLMs are unable to learn or comprehend real-world scenarios involving extremely long contexts, such as understanding entire code repositories or extended video and audio sequences [193]. However, token compression significantly condenses and abstracts original information representations, making it possible for MLLMs to understand these long contexts in a single pass.
Traditional methods for handling long contexts in MLLMs, like FlashAttention [163, 164] or RingAttention [194], involve architectural changes to the model’s attention mechanism to directly accommodate longer sequences. While effective, these require fundamental model modifications. Token compression offers a different, often simpler, route. Instead of redesigning the model to fit more tokens, it focuses on making each token more powerful. By creating information-dense tokens, we pack more meaning into fewer pieces of data. This lets existing MLLM architectures process significantly longer conceptual contexts without major overhauls. It’s a more efficient and accessible way to achieve that crucial one-shot understanding of vast, complex real-world information [195].
@@ 6.4.2 Task-Specific Challenges
AuroraCap [45] Video Caption VDCscore Describe the video in detail.
@@ 6.6.1 Unified Multimodal Token Compression
Although different modalities exhibit unique redundancy patterns requiring specialized compression strategies, contemporary MLLMs predominantly focus on joint multimodal inference [10], whereas unimodal scenarios remain constrained by single-modality inputs. As established in Sections 3, 4, and 5, fundamental algorithmic principles (including transformation-based, similarity-based, attention-based, and query-based approaches) demonstrate transmodal applicability, indicating the viability of developing a unified multimodal token compression framework. Furthermore, leveraging cross-modal correlations represents a promising research direction for future token compression.